OpenStudy (anonymous):

How can we make the number 99 by using any mathematical operation with the condition that we have to use 2 numbers from 1 to 9 for only 1 time.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (watchmath):

You mean we can only take two numbers like 3,4 ?

jhonyy9 (jhonyy9):

like this for example : 9x9=81+2x9=99

jhonyy9 (jhonyy9):

pardon this is correct (4+5)*2+2x9=81+18=99